### v1

# Explosive synchronization in networks of Type-I neurons with electrical synapses

Abstract
Explosive synchronization (ES), which was observed in the scale-free network of the Kuramoto model Gómez-Gardeñes et al. (2011), has been studied widely in the oscillator model. However, investigations of ES in neuronal networks, in spite of their importance in neuroscience, are limited and restricted to specific models. In this work, we explore the nature of the transition to synchronization in a class of neurons, namely Type-I neurons. Leveraging the mapping between Type-I neurons and the Kuramoto model Clusella, Pietras, and Montbrió (2022), we investigate whether the conditions known to induce ES in the Kuramoto model also do so in Type-I neurons. The neurons are coupled through electrical synapses and placed on a scale-free and star networks with complete and partial degree-frequency correlation conditions. Our simulations show ES in networks of Quadratic Integrate and Fire (QIF) neurons, the normal form of Type-I neurons, under weak heterogeneity. We further confirm this phenomenon in networks of Morris-Lecar neurons, in the regime of Type-I excitability, under similar conditions to the QIF neurons. Thus, this work establishes a set of universal conditions that allows ES to arise in Type-I neurons.
preprint: AIP/123-QED
Explosive or abrupt transitions to synchronization have implications for several brain processes, such as epileptic seizure and switching between conscious states. Such transitions have been studied extensively in the abstract Kuramoto model. Although previous studies have reported explosive synchronization (ES) in neuronal networks, they are model-dependent and lack the generality to extend to other neuron models. Our work bridges this gap by demonstrating ES in a class of neurons. We achieve this by establishing ES in the normal form of Type-I neurons — the QIF model, and further affirming it in a Biophysical model exhibiting Type-I behavior. Leveraging the mapping between the Type-I neurons and the Kuramoto model Clusella, Pietras, and Montbrió (2022), we realize theoretical predictions of ES under different conditions on neuronal networks. The generality of these results across a class of neurons, together with their origin in the Kuramoto framework, takes us a step closer to understanding abrupt transitions in the brain.
I Introduction
Synchronization is a widely observed phenomenon across complex systems. Physical systems, such as coupled metronomes Pantaleone (2002), Josephson junctions Wiesenfeld, Colet, and Strogatz (1996), power-grid networks Motter et al. (2013), fireflies Mirollo and Strogatz (1990); McCrea, Ermentrout, and Rubin (2022), and neuronal networks Ward (2003), all show varying degrees of synchronization. The onset of synchronization can occur smoothly, resembling a second order phase transition, or abruptly with accompanying hysteresis as in a first order transition. This abrupt transition is called explosive synchronization (ES). Abrupt synchronization of neurons is hypothesized as the mechanism behind epileptic seizures Jiruska et al. (2013) and also for the onset of Anesthetic-induced unconsciousness Kim et al. (2016). Recent works hypothesize that the brain operates in a critical state close to a phase transition Fontenele et al. (2019); di Santo et al. (2018); Khoshkhou and Montakhab (2019). This emphasizes the need for a better understanding of synchronization in neuronal networks.
Explosive synchronization was first demonstrated in a scale-free network of Kuramoto oscillators with complete degree-frequency correlation Gómez-Gardeñes et al. (2011). Since then, there have been significant efforts in generalizing the conditions for ES in networks of Kuramoto oscillators and their variants. Relaxing the stringency of this first work, imposing only partial correlation Pinto and Saa (2015) and adding disorder to the frequencies Skardal and Arenas (2014) also exhibited ES on scale-free networks. Variants of the Kuramoto model with a phase lag can undergo ES on other network topologies if some extent of degree-frequency correlation exists Kundu and Pal (2019); Kundu et al. (2017). Other studies employing the Kuramoto model have also examined alternate interplays between structures and dynamics, which include frequency weighted coupling Zhang et al. (2013), adaptive coupling Zhang et al. (2015); Avalos-Gaytán et al. (2018), and frequency-gap networks Leyva et al. (2013). Multilayer networks of Kuramoto-Sakaguchi oscillators display ES for a wider class of network topologies, even without any structure-dynamics correlation Kachhvah and Jalan (2017, 2021). Beyond the Kuramoto model, degree-frequency correlated networks of Rössler oscillators and a few experimental systems - Electrical circuits Leyva et al. (2012), mercury beating-oscillator Kumar et al. (2015), and photochemical oscillators Călugăru et al. (2020) - show a sharp transition to synchronization. In several models, it is found that the nature of coupling is a key ingredient in inducing explosive synchronization Sharma (2019, 2021); Ramesan, Shajan, and Shrimali (2022). A more comprehensive account of the existing work on explosive synchronization can be found in this review Bayani, Jafari, and Azarnoush (2022).
Despite the significance of ES to better understand brain processes, studies showing ES in neuronal networks are sparse. The neuron models studied exhibit ES under specific network topologies and with additional model-specific constraints. Izhikevich neurons and Chialvo neurons undergo explosive synchronization on a scale-free network without degree frequency correlation. Izhikevich neurons also exhibit such a sharp transition in the ER network Khoshkhou and Montakhab (2018); Roy, Poria, and Hens (2021). Small-world networks of Hodgkin-Huxley and Chialvo neurons also undergo an explosive synchronization Khoshkhou and Montakhab (2020); Boaretto et al. (2019). The FHN model shows ES with nonlocal coupling Verma and Ambika (2022) and on a degree-frequency correlated scale-free network in the presence of noise Chen et al. (2013). The nature of the transition - first order or second order - in neuronal networks can also depend on the type of synapses connecting the neurons Khoshkhou and Montakhab (2018, 2020). The brain consists of a large variety of neurons, each requiring a specific model. The nature of transitions in these neuronal networks is model specific, limiting a broad understanding of synchronization transitions in the Brain network. To overcome this barrier, we explore explosive synchronization in a class of neurons.
For this work we chose Type-I neurons, a class that is characterized by arbitrarily low frequencies near the onset of oscillations. Type-I neurons are predominantly found in the cortex and hippocampus Prescott, De Koninck, and Sejnowski (2008); Laing et al. (2003). Transition from rest to oscillations in these neurons occurs through the Saddle Node on Invariant Circle bifurcation. The Quadratic Integrate-Fire (QIF) model represents the normal form of all Type-I neurons close to the bifurcation point. Taking the QIF model as a prototype of all Type-I neurons, we explore explosive synchronization transition in networks of QIF neurons. To examine the generality of the results obtained, we further analyze synchronization transitions for networks of Morris-Lecar (ML) neurons, known to display Type-I dynamics Tsumoto et al. (2006). This is a biophysical model originally developed to model the dynamics of barnacle muscle fiber Morris and Lecar (1981).
The networks of QIF neurons with weak heterogeneity and weak electrical coupling map to the standard Kuramoto model Clusella, Pietras, and Montbrió (2022). Building on the well documented studies of explosive synchronization in the Kuramoto oscillator network under various conditions, we explore the emergence of explosive synchronization in Type-I neuronal networks. Brain connectivity is known to exhibit a scale-free topologyVan Den Heuvel and Sporns (2013); van den Heuvel and Sporns (2011). Numerous studies with the Kuramoto model on scale free networks have shown explosive Synchronization under different conditions. We leverage the QIF-Kuramoto model mapping to study the synchronization transition of scale-free networks of Type-I neurons with varying degrees of correlation between structure and dynamics.
This paper is organized as follows. In section II we describe the neuron models used in this work with details of the models, commenting on the nature of the bifurcations in these models. In sections III.1 and III.2, results of numerical simulations carried out for networks of QIF neurons and ML neurons are presented, respectively. Section IV gives the concluding remarks of the results discussed in the previous sections.
II Model
In Type-I neurons, the change from excitatory to oscillatory state is achieved by the injected current and occurs through a saddle Node on invariant circle (SNIC) bifurcation. The Quadratic Integrate-and-Fire (QIF) model is a simple yet powerful neuronal model that captures the essential dynamics of Type-I neurons. The following equations describe the QIF model:
$\tau\dot{V_{i}}=V_{i}^{2}+\eta_{i}+\varepsilon I_{i,syn}(t)$ (1a)
$if\ V_{i}>V_{p},\ then\ \ V_{i}\rightarrow V_{r}$ (1b)
where $V_{i}$ is the voltage of $i^{th}$ neuron in the network. The term $\eta_{i}$ represents the external current injected into the neuron, and the term $I_{i,syn}$ represents the net synaptic current flowing into neuron i. When $V_{i}$ reaches a large threshold value, $V_{p}$ (often approximated as $\infty$), it is reset to a large negative value, $V_{r}$ (often taken as -$\infty$), to mimic the action potential firing and reset process of a neuron. Eq. (1a) describes the approach of a QIF neuron to the threshold by the normal form of SNIC bifurcation, while Eq. (1b) represents the reset process. When $\eta_{i}<0$, the neuron is in an excitable state, and if $\eta_{i}>0$, it oscillates, mimicking a train of action potentials (spikes). The frequency of spiking increases gradually with $\eta_{i}>0$ from an arbitrarily low value - a characteristic of Type-I neurons. A well known variable transformed version of the QIF model is the $\Theta$-neuron model" Ermentrout and Kopell (1986). In a recent work Clusella, Pietras, and Montbrió (2022), the QIF model was shown to reduce to the Kuramoto-Sakaguchi model with $V_{p}=-V_{r}=\infty$, under weak synaptic coupling and weak heterogeneity.
In our work, we couple the QIF neurons through electrical synapses, leading to the following form for Eq. (1a).
$\tau\dot{V_{i}}=V_{i}^{2}+\eta_{i}+\varepsilon g\sum_{j=1}^{N}A_{ij}(V_{j}-V_{i})$ (2)
while the reset condition remains the same. In the above equation, g is the strength of the electrical synapse, $A$ is the adjacency matrix which characterizes the network topology with $A_{ij}=1$ if $i^{th}$ neuron is connected to the $j^{th}$ neuron, and $A_{ij}=0$ if there is no connection between them. There exists no self-connection, i.e., $A_{ii}=0$ for all $i$. We consider two network topologies - scale-free (SF) and star networks. The SF network has a degree distribution given by the power law $p(k)=k^{-\gamma}$ where the degree $k_{i}=\sum_{j}^{N}A_{ij}$ of neuron $i$ is the number of neurons in the network connected to it. In a star network, the single hub neuron is connected to all the other neurons, which are not themselves interconnected. An SF network is characterized by hub-like structures. Hence, star networks act as a precursor to the SF network. We consider networks of QIF neurons with varying degrees of correlation between their degrees and the frequencies. Since for a QIF neuron, $\eta$ is proportional to its frequencyErmentrout and Kopell (1986), degree-frequency correlation is imposed following the rule $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$.
Between two successive spikes, voltage dynamics of a free-running QIF neuron is governed by Eq. (2) with $g=0$. Its solution Pfeuty et al. (2003), starting from the voltage value $V_{r}$ at $t=0$ is
$V_{i}(t)=\sqrt{\eta_{i}}\tan\Bigg[\frac{t\sqrt{\eta_{i}}}{\tau}+\arctan\Bigg(\frac{V_{r}}{\sqrt{\eta_{i}}}\Bigg)\Bigg]$ (3)
with its firing frequency $\Omega_{i}={(\pi\sqrt{\eta_{i}})}/{(\tau\arctan({V_{p}}/{\sqrt{\eta_{i}}}))}$. The phase $\theta_{i}=\Omega_{i}\ t$ of a QIF neuron can be expressed in terms of its voltage as
$\theta_{i}=\bigg[\frac{\pi}{\arctan(\frac{V_{p}}{\sqrt{\eta_{i}}})})\bigg]\bigg[\arctan(\frac{V_{i}}{\sqrt{\eta_{i}}})-\arctan(\frac{V_{r}}{\sqrt{\eta_{i}}})\bigg]$ (4)
Accordingly, $\theta_{i}\rightarrow 2\pi$ as $V_{i}(t)\rightarrow V_{p}$ and $\theta_{i}\rightarrow 0$ as $V_{i}(t)\rightarrow V_{r}$. To quantify the degree of synchronization, we use an instantaneous order parameter $r(t)=\frac{1}{N}|\sum_{j=1}^{N}e^{i\theta_{j}(t)}|$, which is time averaged over steady states to yield a global order parameter $R=\langle r(t)\rangle$. R varies from 0 to $1$, with $R=1$ representing the completely synchronized state and $R=0$ the asynchronous state of the network.
The Morris-Lecar model is a conductance based biophysical model formulated to describe the dynamics of the giant barnacle muscle fibre. The following equations constitute the dynamics of a Morris-Lecar neuron connected to other neurons in the network through electrical synapses.
$C_{M}\frac{dV_{i}}{dt}=$ $-{g_{L}}(V_{i}-V_{l})-g_{Ca}M_{\infty}(V_{i}-V_{Ca})-g_{K}W_{i}(V_{i}-V_{K})$ (5a)
$+I_{i,ext}-g_{syn}\sum_{j=1}^{N}A_{ij}(V_{j}-V_{i})$
$\frac{dW_{i}}{dt}=$ $\frac{W_{\infty}-W_{i}}{\tau_{W}}$ (5b)
In Eq. (5b), $W_{i}$ represents the recovery variable of the $i^{th}$ neuron. $V_{l}$, $V_{Ca}$, $V_{K}$ represents the equilibrium potentials of the leak, $Ca^{2+}$, and $K^{+}$ currents respectively. $g_{l}$, $g_{Ca}$, and $g_{K}$ are the maximum conductance of the corresponding ionic channels. In our work, we have chosen $V_{l}=-60mV$, $V_{Ca}=120mV$, $V_{K}=-80mV$, $g_{l}=2\ mS/cm^{2}$, $g_{Ca}=2\ mS/cm^{2}$, and $g_{K}=4\ mS/cm^{2}$. The capacitance of the neuron membrane, $C_{M}$ is taken as $20\mu F/Cm^{2}$. $M_{\infty}(V_{i})=0.5[1+\tanh{\{(V_{i}-V_{1})/V_{2}\}}]$ and $W_{\infty}(V_{i})=0.5[1+\tanh{\{(V_{i}-V_{3})/V_{4}}\}]$ are the steady-state activation function of the $Ca^{2+}$ and $K^{+}$ currents respectively and $\tau_{w}(V_{i})=1/[\phi\cosh{\{(V_{i}-V_{3})/2V_{4}\}}]$ is the time constant of the $K^{+}$ activation, with $V_{1}=-1.2mV$, $V_{2}=18.0mV$, $V_{3}=12.0mV$, and $V_{4}=17.4mV$. These parameter values, chosen from Tsumoto et al. (2006), ensure that the neurons are operating in the Type-I regime. External current, $I_{i,ext}$, drives the neurons from an excitable to an oscillatory state through SNIC bifurcation. In an oscillatory state, the free-running frequency of an uncoupled neuron increases with the external current. We consider the same network topologies as those for QIF neurons. Degree-frequency correlation is imposed by fixing $I_{i,ext}=I_{0}+\varepsilon k_{i}$. The value of $I_{0}$ is chosen to keep the ML neuron close to the bifurcation point so that its dynamics can be approximated by the QIF model. The phase of a neuron’s dynamics is expressed by
$\phi_{i}(t)=2\pi\frac{t-T_{n,i}}{T_{n+1,i}-T_{n,i}},\ \ T_{n,i}\leq t<T_{n+1,i}$ (6)
where $T_{n,i}$ and $T_{n+1,i}$ are the $n^{th}$ and $(n+1)^{th}$ spike times respectively of the $i^{th}$ neuron. These phases are used to calculate the instantaneous order parameter and thereby the global order parameter as described earlier.
III Results
In this section, we discuss the results of numerical simulations carried out for various networks of QIF and ML neurons connected through electrical synapses.
III.1 Networks of QIF neurons
We simulated 1000 QIF neurons arranged on SF networks built via a configuration model. This model gives control over the degree exponent $\gamma$ of the degree distribution $p(k)$. We have considered $2<\gamma<3$ in this work. The dynamics of QIF neurons are governed by Eq. (2) along with the reset rule of Eq. (1b). We fix $V_{p}=-V_{r}=750$ and $\tau=1$ for all neurons. Synaptic strength $g$ is identical between all connected neurons. Simulations are performed into steady states for increasing values of $g$, starting at $g=0$, to a maximum $g$ value. Thereafter, simulations are similarly carried out for decreasing values of $g$. Final steady state voltages for the current $g$ are used as the initial conditions for the next $g$ value. At $g=0$, we start with uniformly distributed voltage values. The order parameter is calculated in the steady state for every value of $g$ in forward and backward runs. We solve Eq. (2) using Euler method with a small step size dtdt ($\approx 10^{-4}$) over the duration $T_{f}=10000$. First $5000$ time units are discarded as transients, and the rest are used for evaluating $R$.
Figure 1: A sharp transition to synchronization accompanied by hysteresis in the order parameter is seen for a scale-free network of 1000 QIF neurons with complete degree-frequency correlation. $\gamma=2.4$ (black line) and $\gamma=2.7$ (red line). For both cases $\bar{\eta}=20$, $\varepsilon=0.01$ with solid curves representing forward simulations and dotted lines representing backward simulations.
Figure 2: Plots of 1000 QIF neurons connected by an SF network with varying degree-frequency correlations. In all cases, the same SF network with $\gamma=2.3$ is used. (a) Complete degree-frequency correlation. (b) Only the top $10\%$ of neurons with the highest degree are assigned frequencies correlated with their degree, while the rest are assigned random frequencies. (c) All neurons are assigned random frequencies. In both (b) and (c), random frequencies are drawn from a Lorentzian distribution $p(k)=a/\pi((k-k_{0})^{2}-a^{2})$ with $k_{0}=0$ and $a=1$. For all cases $\bar{\eta}=20$ and $\varepsilon=0.01$.
A complete degree-frequency correlation is imposed by setting $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ for every neuron. We choose $\bar{\eta}>0$ to place neurons in the oscillatory regime. To maintain weak heterogeneity in $\eta$ values across neurons, we take $\bar{\eta}>>\varepsilon$, with $\bar{\eta}=20$ and $\varepsilon\approx 10^{-2}\in(0.005,0.01)$. With a small $\varepsilon$, the coupling is weak between neurons. Variations in the order parameter while increasing and decreasing values of $g$ are shown in Fig. 1. A sharp transition to synchronization along with hysteresis is clearly seen - indicative of an explosive synchronization. This result is robust to the degree exponent of the scale-free network, as shown in fig (1) for different $\gamma$ values.
wExploiting the mapping between the QIF neuronal network and the Kuramoto model Clusella, Pietras, and Montbrió (2022), we also investigated synchronization transition in partially degree-frequency correlated QIF networks, which are known to show ES in the Kuramoto networks Pinto and Saa (2015). To achieve partial degree-frequency correlation in the QIF network, a fraction of the neuron population is assigned frequencies following the $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ rule, while the rest are assigned random frequencies. Fig. 2 shows that the first-order transition to synchronization observed in a fully degree-frequency correlated SF network of QIF neurons persists even when only $10\%$ of the highest degree neurons are correlated. In contrast, when the frequencies of all the neurons are uncorrelated with their degrees, the transition becomes continuous.
Next, we describe transitions in star networks of fully degree-frequency correlated QIF neurons. We assign frequencies following $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ for every neuron. Simulations are performed by varying g following the same method as for the scale-free networks. Eq. (2) is solved using the RK4 method with step size dtdt ($\approx 10^{-4}$). Simulations are carried out till $T_{f}=25000$, and the order parameter is calculated over the final 10000 time units. Fig. 3 shows a sharp transition to synchronization across all network sizes. The forward transition points shift to lower values of $g$ as the size of the star network is increased.
Figure 3: Forward simulations of Star networks of QIF neurons with complete degree-frequency correlation. N is the number of neurons in the network. For these simulations $\bar{\eta}=20$, $\varepsilon=0.01$.
Figure 4: Effect of $\varepsilon$ on the transition points in a 21 complete degree-frequency correlated star networks of QIF neurons. The solid lines represent the forward g simulation, while the dotted line represents the backward simulation. Other parameters are the same as in Fig. 3.
A network of QIF neurons with pure electrical coupling is known to reduce to the standard Kuramoto model Clusella, Pietras, and Montbrió (2022), yielding the following expression for its phase evolution.
$\dot{\theta_{i}}=\bigg(\frac{2\sqrt{\bar{\eta}}}{\tau}+\frac{\varepsilon k_{i}}{\tau\sqrt{\bar{\eta}}}\bigg)+\frac{\varepsilon g}{\tau}\sum_{j}^{N}\sin(\theta_{j}-\theta_{i})$ (7)
where $\omega_{i}=(\frac{2\sqrt{\bar{\eta}}}{\tau}+\frac{\varepsilon k_{i}}{\tau\sqrt{\bar{\eta}}})$ and $\lambda=\varepsilon g/\tau$ are the frequency and the coupling strength respectively of the reduced Kuramoto model. In Fig. 4, the results of simulating a complete degree-frequency correlated star network of 21 QIF neurons for different levels of heterogeneity characterized by $\varepsilon$ values are shown. In these simulations, when $g$ values are less than a certain value Leyva et al. (2025), a small randomness is added to the $\eta$ values during the backward run. This is done to slightly perturb the system out of the synchronized branch. These plots show that increasing $\varepsilon$ shifts the forward transition point ($\lambda_{cf}$) to a higher value. This trend aligns with the increase in the forward transition point with the rise in heterogeneity as reported for the Kuramoto network in W. Liu et al Liu et al. (2013). For a complete degree-frequency correlated star network with K leaves, the hub frequency is $\omega_{h}=K$, and the frequency of the leaves is $\omega_{l}=1$. The mean-field calculation carried out for such a network of Kuramoto oscillators gives the critical coupling strength $\lambda_{c}=(\Omega-\omega_{l})=\frac{K-1}{K+1}$ where $\Omega$ is the average frequency of the network Gómez-Gardeñes et al. (2011). Below this, a stable synchronized state does not exist. From the expression of $\omega_{i}$ in the Eq. (7), the average frequency of the reduced Kuramoto model will be $\Omega=(2\pi/\tau)+(2\varepsilon K/(\tau\bar{\eta}(K+1)))$. Thus, the corresponding theoretical $\lambda_{c}$ for the star network of QIF neurons is
$\lambda_{c}=\frac{\varepsilon g_{c}}{\tau}=\frac{\varepsilon(K-1)}{\tau\bar{\eta}(K+1)}$ (8)
For $\varepsilon=0.0095$ and $\varepsilon=0.0065$, theoretically predicted $\lambda_{c}$ values are $\approx 0.0019$ and $\approx 0.0013$ respectively. In Fig. (4), it is seen that the backward transitions occur at $\lambda$’s close to these values predicted by the above equation.
III.2 Networks of Morris-Lecar neurons
In the previous section, we reported that the SF networks of the normal form of Type-I neurons (the QIF model) show explosive synchronization with complete and partial degree-frequency correlation. However, the transition becomes smooth if the degrees and the frequencies are uncorrelated. Motivated by these results, we consider Type-I Morris-Lecar neurons on the same network topologies with different levels of degree-frequency correlation. A Morris-Lecar neuron, modeled by Eq. (5), shows Type-I behavior for the chosen parameter values of section II. With these parameter values a SNIC bifurcation, separating excitable and oscillatory states, occurs at $I_{ext}\approx 39.6$ Tsumoto et al. (2006). Following section II, we correlate the frequency of a neuron with its degree as per the rule $I_{i,ext}=I_{0}+\varepsilon k_{i}$, with $I_{0}>39.6$. Simulations were performed for scale-free networks of 1000 ML neurons. The strength of the electrical synapse was varied following the procedure considered for scale-free networks of QIF neurons. The order parameter was calculated using the phases defined in Eq. (6).
First, we consider complete degree-frequency correlation with different $\gamma$ values of the scale-free network. Results are shown in Fig. 5. The transition to synchronization is explosive with a sharp jump in the order parameter accompanied by hysteresis.
[…]
Figure 6: Plots of 1000 ML neurons connected through an SF network. In all cases, the same SF network with $\gamma=2.4$ is used. (a) Complete degree-frequency correlation. (b) Only the top $10\%$ of neurons with the highest degree are assigned frequencies correlated with their degree, while the rest are assigned random frequencies. (c) All neurons are assigned random frequencies. In both (b) and (c), random frequencies are drawn from a Lorentzian distribution $p(k)=a/\pi((k-k_{0})^{2}-a^{2})$ with $k_{0}=0$ and $a=1$. For all cases $I_{0}=43.0$ and $\varepsilon=0.01$.
Partial degree-frequency correlation is achieved by assigning frequencies to the uncorrelated neurons in the same way as was done for the QIF networks. Simulation results are shown in Fig. 6. We observe an abrupt transition to a synchronized state with only $10\%$ of correlation. When the neuron frequencies are completely uncorrelated to their degrees, the transition becomes continuous.
With these results of partial degree-frequency correlation simulations of ML neuronal network, we show that any Type-I neurons, with weak heterogeneity and coupling, can show ES if the neuronal network satisfies the conditions needed for ES in the corresponding Kuramoto model. These observations verify the mapping among these three models, which approximates the ML model with the QIF model, and the QIF model with the Kuramoto model.
IV Conclusion
We have demonstrated the emergence of explosive synchronization in scale-free and star networks of Type-I neurons with electrical coupling and weak intrinsic heterogeneity. Our findings broaden the presence of ES, showing that it is not limited to specific neuronal models but can arise across a broad class of neurons. We establish this by demonstrating ES in the normal form of Type-I neurons (QIF model) and affirming it in a biophysical model of Type-I neurons (Morris-Lecar model). Our results in Fig. 1, 2 ,5 ,6 show that ES is robust to the degree exponent of scale-free networks and to various levels of degree-frequency correlations. For ES on star networks of QIF neurons, there was quantitative agreement between simulations and theory. Specifically, the critical coupling for backward transitions in numerical simulation agrees with theoretical predictions from the mapped Kuramoto model.
ES is hypothesized as the mechanism for epileptic seizures and the onset of anesthetic-induced unconsciousness. Understanding the conditions of explosive synchronization in neuronal networks could shed light on normal and pathological brain dynamics. Leveraging the mapping between Type-I neurons and the Kuramoto model, a large set of scenarios can be explored for explosive synchronization. This work is the first step towards it.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.